# Towards Small Language Models for Security Query Generation in SOC Workflows

### v1

@@ 1. Introduction
The evolving cybersecurity landscape has increased attack complexity and reshaped how analysts mitigate threats. In Security Operations Centers (SOCs), analysts face overwhelming volumes of event logs, network traffic, and threat intelligence feeds (Bargan_2024), receiving on average 5,000 alerts per day via SIEM systems and up to 100,000 in extreme cases (Alahmadi). To investigate these massive logs, they rely on query languages such as Sigma, Elastic EQL, and KQL (SigmaHQ; Elastic_Query_Language_Documentation; Kusto-Query-Language), but translating natural-language intent into correct queries over large, evolving schemas remains a major bottleneck.
KQL, introduced by Microsoft in 2017 (KQL_Background), is the most widely adopted. Integrated into Microsoft Sentinel and Defender, it is a domain-specific, read-only language that lets analysts parse millions of log rows efficiently and extract relevant events. However, its expressiveness and non-trivial semantics make authoring accurate queries challenging for non-experts, especially during time-sensitive investigations.
In the era of LLMs, automated KQL generation can further accelerate investigations by producing precise queries aligned with an analyst’s intent. Beyond code generation, LLMs have been applied across cybersecurity tasks, such as log analysis and penetration testing (PentestGPT; VulDetect). They have also been exploited for offensive purposes, including phishing and ransomware planning (Hassanin_Moustafa_2024), while models like GPT-3.5 and GPT-4 show strong defensive potential in areas such as scam detection (Jiang_2024). However, LLM-centric pipelines for NLQ-to-KQL translation can be costly, latency-sensitive, and difficult to deploy under enterprise data-governance constraints.
@@
Unfortunately, LLMs also incur drawbacks, including higher latency, memory demands, and operational costs. Moreover, depending on the use case, they may be unsuitable for certain tasks (Belcak_Heinrich_Diao_Fu_Dong_Muralidharan_Lin_Molchanov). As an alternative, we introduce SLMs, defined following (Belcak_Heinrich_Diao_Fu_Dong_Muralidharan_Lin_Molchanov) as language models that (i) can run on common consumer devices with practical inference latency, (ii) are not LLMs, and (iii) have at most 10 billion parameters. SLMs offer lower latency, memory, computational, and operational costs while still achieving accuracy comparable to LLMs in some domain-specific tasks (Belcak_Heinrich_Diao_Fu_Dong_Muralidharan_Lin_Molchanov). Despite this promise, there has been no systematic study of SLMs for NLQ-to-KQL translation in realistic, schema-rich settings.
@@
We further explore LoRA fine-tuning (LoRA) with rationale distillation to transfer reasoning from a teacher LLM into an SLM. LoRA trains only low-rank adapter parameters while freezing the original model weights, enabling efficient fine-tuning on NLQ-KQL pairs. To embed reasoning capabilities, each training example is augmented with a short chain-of-thought explanation produced by the teacher model, followed by the target KQL. This rationale-augmented setup allows the SLM to learn intermediate reasoning steps in addition to final outputs, strengthening its ability to handle structured code generation tasks without increasing model size.
@@
We conduct extensive experiments spanning baseline performance, prompting strategies, fine-tuning, and architectural design. Our evaluation covers nine models: five SLMs (Gemma-3-1B-IT, Gemma-3-4B-IT, Phi-4-Mini-Instruct,Qwen-2.5-7B-Instruct-1M, and DeepSeek Coder 6.7B Instruct) and four LLMs (GPT-5, GPT-4o, Gemini 2.0 Flash, and Phi-4), using 230 NLQ-KQL pairs from Microsoft’s NL2KQL Defender Evaluation Dataset (NL2KQL_Eval) and 83 additional queries from Sentinel-Queries (reprise99) for generalization. LLMs achieve high syntactic accuracy (above 0.90) but show variable semantic correctness, while SLMs perform poorly in zero-shot settings, with semantic scores near 0 and table/filter scores below 0.1. Targeted prompting and LoRA fine-tuning with rationale distillation significantly improve SLM performance, especially for DeepSeek Coder 6.7B Instruct with NL2KQL. Our two-stage SLM-Oracle architecture achieves a syntax score of 0.987 and semantic score of 0.906 on Microsoft’s NL2KQL Defender Evaluation dataset. We demonstrate generalizability on Microsoft Sentinel data with a syntax score of 0.964 and semantic score of 0.831, approaching LLM performance while remaining up to 10x cheaper in token cost than GPT-5. These findings show that SLMs, when combined with lightweight LLM refinement, enable accurate and cost-efficient KQL generation at scale.
@@ 2. Related Work
Natural Language to Query Generation Converting NLQs to structured queries has been widely studied across domains and languages. Early work (popescu2003towards; li2014nalir; baik2019bridging) translated NLQs to SQL using rule-based methods and custom heuristics, but extending these techniques to new databases and languages demands substantial manual effort and struggles with edge cases (nl2kql). Neural methods improve portability by casting text-to-SQL as sequence generation with integrated validation. PICARD (PICARD) uses sequence-to-sequence models with modular syntactic checks to ensure query correctness and alignment with the NLQ. Supplying schema context further improves semantic fidelity (Hong_Yuan_Chen_Zhang_Huang_Huang_2024). In practice, these ideas combine constrained decoding for syntax with schema-aware prompting or retrieval for semantics, directly informing our KQL setting. Unlike prior systems that center on a single model, our approach uses both SLMs and LLMs. These general principles have been extended to domain-specific query languages, especially in security, where KQL is the primary language for threat hunting and log analysis. NL2KQL (nl2kql) translates NLQs to KQL by combining embedding-based semantic similarity with few-shot prompting in a modular, end-to-end framework. XPert (X_Pert) takes an incident-centric approach, retrieving similar historical tickets to recommend or compose queries using in-context learning with embedding-based retrieval. However, both systems rely heavily on LLMs, resulting in cost and latency constraints that limit real-time SOC deployment, where analysts need rapid query assistance. In contrast, our approach uses an SLM for generation, avoiding exclusive dependence on LLMs.
@@
Security-Specific LLMs LLMs are increasingly applied across SOC tasks and offensive evaluations, but their effectiveness hinges on task-specific scaffolding and verification. VulDetect (VulDetect) fine-tunes GPT-style models for anomaly detection in logs using a transformer-based framework. PentestGPT (PentestGPT) builds an LLM penetration-testing assistant with modular task support, reporting bounded gains on real targets, highlighting the importance of structured tooling around the model. Although LLMs offer strong code generation capabilities, their cost and latency often make real-time deployment impractical (Belcak_Heinrich_Diao_Fu_Dong_Muralidharan_Lin_Molchanov), pushing real-time operations toward SLMs.
@@
SLMs Adoption SLMs provide attractive latency and cost profiles for SOC settings but require augmentation to match LLM quality. SuperICL shows that small, locally fine-tuned models can act as plug-ins to larger LLMs, structuring context while the larger model executes—a useful division of labor for SOC workflows (SLM_Ref). Parameter-efficient fine-tuning further narrows this gap while preserving efficiency; methods like LoRA and QLoRA enable targeted adaptation on modest hardware (LoRA; Improving_Phishing_Email_Detection; qlora). Complementary techniques such as schema/value retrieval, few-shot selection, and best-of-N with an external judge improve correctness without enlarging model footprint (wang2022self; madaan2023self). In sum, SLMs are viable for SOC use when paired with retrieval, schema hints, and parameter-efficient tuning.
@@ 3.2. Zero-Shot Prompting
The Zero-Shot prompting strategy aims to interact with the model without containing any relevant examples or demonstrations (Prompt_Engineering_Guide). Zero-Shot prompting can be an extremely powerful technique, as it typically utilizes less tokens than other prompting strategies and relies mostly on the inference capabilities of the language model itself. Furthermore, instruction tuning has been shown to improve Zero-Shot learning capabilities (Payong_2024).
@@ 3.3. NL2KQL-Inspired Prompting
In addition to Zero-Shot prompting, we re-implement NL2KQL (nl2kql), the state-of-the-art system for KQL query generation. NL2KQL consists of five main components: Semantic Data Catalog, Schema Refiner, Few-Shot Selector, Prompt Builder, and Query Refiner. For detailed descriptions of these components, we refer the reader to the original paper (nl2kql). Below, we describe the changes we made in our recreation.
@@ 3.3.2. Schema Refiner
Consistent with (nl2kql), the Schema Refiner retrieves the top-$t$ relevant tables and associated columns using cosine similarity between the NLQ and the embeddings. As in the original, $t=9$ and $v_{\text{n}}=5$. We follow the same procedure but base our embeddings on the recreated Semantic Data Catalog.
@@ 3.3.4. Prompt Builder
We preserve the design of the Prompt Builder from (nl2kql), which combines detailed instructions, syntax rules, best practices, and few-shot examples into a single prompt template.
@@ 3.3.5. Query Refiner
We implement the Query Refiner following (nl2kql), which uses Microsoft’s KQL Parser (Kusto-Query-Language) to verify syntactic and semantic correctness, detect undefined identifiers, repair aggregate functions, fix parentheses, and add missing operators. We retain the embedding-based replacement strategy with a cosine similarity threshold of 0.9.
@@ 3.3.6. Alternative Prompting Strategies
Beyond faithfully reimplementing NL2KQL (nl2kql), we introduce prompting strategies that supply models with targeted tips on common errors. The first focuses on frequent syntactic issues identified via the KQL Parser. The second adds guidance for semantic errors such as mismatched column types and invalid table references. We compare both strategies and adopt the one with the highest correctness. Due to SLMs’ limited inference capabilities, concise, model-specific tips help them generate more accurate KQL compared to supplying the full syntax and semantic rules. Our reimplementation substitutes Google’s text-embedding-004 for embeddings, uses Google Gemini 2.0 Flash and Microsoft Phi-4 for query generation, and shifts correctness filtering to post-generation. We further extend prompting with error-aware instructions to assess whether targeted guidance improves KQL generation.
@@ 3.4. LoRA Fine-Tuning with LLM Distillation
To further assess the quality of KQL queries that are generated from SLMs, we also test LoRA, a fine-tuning method that freezes pre-trained model weights from language models, and applies decompositional matrices to transformers layers within the LLM in order to alter a subset of model weights (LoRA). The goal of this method is to train SLMs on NLQ-KQL pairs so that SLMs can be further improved in producing quality responses. LoRA can be mathematically defined as follows:
@@
Where B $\in$ $\mathbb{R}^{d\times r}$, A $\in$ $\mathbb{R}^{r\times k}$, and both are low-rank matrices. During training time, the values of the original weight matrix $\mathcal{W}$ remain stable while $\Delta{W}$ (ABAB) is updated. Altering low-rank matrices instead of the entire weight matrix reduces the memory requirements, gradient storage, and optimizer states needed to perform fine-tuning (Improving_Phishing_Email_Detection). To train the DeepSeek Coder 6.7B Instruct, we use a synthesized dataset created through LLM Knowledge Distillation.
LLM Knowledge Distillation is a transfer learning technique in Machine Learning that is used to relay the inference and reasoning capabilities of larger, proprietary LLMs to smaller SLMs. This allows the SLMs to preserve its nature of greater efficiency, deployment feasibility, and faster inference capabilities while also receiving the knowledge of LLMs in its own reasoning processes (Distillation).
@@ 3.5. Two-Staged Architecture
The oracle operates in two modes. In refinement-only, it selects and edits the best SLM candidate using the judge model’s internal knowledge. In schema-aware refinement, it performs the same process but with explicit schema context. We evaluate both modes using two prompts (Figure 2). The oracle is instantiated as an LLM-as-a-Judge, leveraging evidence that judge models align well with human preferences while exhibiting recognizable biases (zheng2023judging).
@@ 4. Evaluation
We evaluate the effectiveness of our proposed design through a series of experiments conducted on a machine with an AMD EPYC 9534 64-Core Processor, an NVIDIA H100 NVL GPU, and Ubuntu 22.04.5 LTS. We address seven research questions (RQs) as follows. For all RQs, we access and query SLMs using the Huggingface Python package (Huggingface), which avoids input/output token rate limits. For RQ5 and RQ6, we query Google Gemini 2.0 Flash via the Google GenAI package (PyPI). Our re-implemented NL2KQL pipeline (Section 3.3) is evaluated using Gemini 2.0 Flash, Phi-4, and GPT-4o.
@@
Datasets RQ1-RQ5 are evaluated using Microsoft’s NL2KQL Defender Evaluation Dataset (NL2KQL_Eval), which contains 230 NLQ-KQL pairs referencing all tables in the NL2KQL Defender Schema. Instead of feeding NLQs directly to the models, we embed them in different prompting strategies and compare the generated queries against the baseline pairs to measure syntactic, semantic, and structural similarity. For RQ6, we obtained 83 NLQ-KQL pairs from the Microsoft Defender sections of Sentinel-Queries (reprise99), a publicly available GitHub repository. We used all available Defender queries from this repository and excluded queries designed for other data sources.
Evaluation Metrics In order to evaluate these research questions and assess the effectiveness of our research methods, we first define the following evaluation metrics. These evaluation metrics have been previously defined in NL2KQL (nl2kql), and we adapt them as part of our evaluation process as well.
@@
Cost:. Total cost (USD) to process 230 NLQs from the NL2KQL Defender Evaluation Dataset, accounting for all input and output tokens per model. SLMs offer several advantages, including enhanced data security and lower response latency, but cost efficiency is one of their most significant benefits. Like LLMs, SLM costs are determined by input and output token counts, with per-million-token pricing varying by provider. For this analysis, we report prices based on Cloudflare’s Workers AI pricing, as highlighted in (The_Cloudflare_Blog_2025). Table 1 presents the cost per 1M tokens for multiple LLMs and SLMs in USD. The cost gap is substantial: SLMs can offer up to 7.5x savings on input tokens alone compared to LLMs. When used effectively for code generation, SLMs can deliver significant performance at a fraction of traditional LLM costs.
@@
Embedding Model An embedding model is an ML model that can be used to convert text into a numerical representation so that it can be conceptually understood by LLMs or SLMs. The primary embedding model used in the following RQs is Google’s text-embedding-004 model. The text-embedding-004 provides embedding queries for \$0.025 per 1 million input tokens, and \$0.02 per 1 million output tokens (Google). Furthermore, text-embedding-004 is a lightweight, simplistic embedding model that has performed well across multiple numerous benchmarks and has strong multilingual and domain-specific capabilities.

### v2

@@ 1. Introduction
The evolving cybersecurity landscape has increased attack complexity and reshaped how analysts mitigate threats. In Security Operations Centers (SOCs), analysts face overwhelming volumes of event logs, network traffic, and threat intelligence feeds (Bargan, 2024), receiving on average 5,000 alerts per day via SIEM systems and up to 100,000 in extreme cases (Alahmadi et al., 2022). To investigate these massive logs, they rely on query languages such as Sigma, Elastic EQL, and KQL (SigmaHQ (2017); Elastic (2018); 17), but translating natural-language intent into correct queries over large, evolving schemas remains a major bottleneck.
KQL, introduced by Microsoft in 2017 (Lean, 2025), is the most widely adopted. Integrated into Microsoft Sentinel and Defender, it is a domain-specific, read-only language that lets analysts parse millions of log rows efficiently and extract relevant events. However, its expressiveness and non-trivial semantics make authoring accurate queries challenging for non-experts, especially during time-sensitive investigations.
In the era of LLMs, automated KQL generation can further accelerate investigations by producing precise queries aligned with an analyst’s intent. Beyond code generation, LLMs have been applied across cybersecurity tasks, such as log analysis and penetration testing (Deng et al., 2024; Omar and Shiaeles, 2023). They have also been exploited for offensive purposes, including phishing and ransomware planning (Hassanin and Moustafa, 2024), while models like GPT-3.5 and GPT-4 show strong defensive potential in areas such as scam detection (Jiang, 2024). However, LLM-centric pipelines for NLQ-to-KQL translation can be costly, latency-sensitive, and difficult to deploy under enterprise data-governance constraints.
@@
Unfortunately, LLMs also incur drawbacks, including higher latency, memory demands, and operational costs. Moreover, depending on the use case, they may be unsuitable for certain tasks (Belcak et al., 2025). As an alternative, we introduce SLMs, defined following (Belcak et al., 2025) as language models that (i) can run on common consumer devices with practical inference latency, (ii) are not LLMs, and (iii) have at most 10 billion parameters. SLMs offer lower latency, memory, computational, and operational costs while still achieving accuracy comparable to LLMs in some domain-specific tasks (Belcak et al., 2025). Despite this promise, there has been no systematic study of SLMs for NLQ-to-KQL translation in realistic, schema-rich settings.
@@
We further explore LoRA fine-tuning (Hu et al., 2022) with rationale distillation to transfer reasoning from a teacher LLM into an SLM. LoRA trains only low-rank adapter parameters while freezing the original model weights, enabling efficient fine-tuning on NLQ-KQL pairs. To embed reasoning capabilities, each training example is augmented with a short chain-of-thought explanation produced by the teacher model, followed by the target KQL. This rationale-augmented setup allows the SLM to learn intermediate reasoning steps in addition to final outputs, strengthening its ability to handle structured code generation tasks without increasing model size.
@@
We conduct extensive experiments spanning baseline performance, prompting strategies, fine-tuning, and architectural design. Our evaluation covers nine models: five SLMs (Gemma-3-1B-IT, Gemma-3-4B-IT, Phi-4-Mini-Instruct,Qwen-2.5-7B-Instruct-1M, and DeepSeek Coder 6.7B Instruct) and four LLMs (GPT-5, GPT-4o, Gemini 2.0 Flash, and Phi-4), using 230 NLQ-KQL pairs from Microsoft’s NL2KQL Defender Evaluation Dataset (22) and 83 additional queries from Sentinel-Queries (Zorich, 2021) for generalization. LLMs achieve high syntactic accuracy (above 0.90) but show variable semantic correctness, while SLMs perform poorly in zero-shot settings, with semantic scores near 0 and table/filter scores below 0.1. Targeted prompting and LoRA fine-tuning with rationale distillation significantly improve SLM performance, especially for DeepSeek Coder 6.7B Instruct with NL2KQL. Our two-stage SLM-Oracle architecture achieves a syntax score of 0.987 and semantic score of 0.906 on Microsoft’s NL2KQL Defender Evaluation dataset. We demonstrate generalizability on Microsoft Sentinel data with a syntax score of 0.964 and semantic score of 0.831, approaching LLM performance while remaining up to 10x cheaper in token cost than GPT-5. These findings show that SLMs, when combined with lightweight LLM refinement, enable accurate and cost-efficient KQL generation at scale.
@@ 2. Related Work
Natural Language to Query Generation Converting NLQs to structured queries has been widely studied across domains and languages. Early work (Popescu et al., 2003; Li and Jagadish, 2014; Baik et al., 2019) translated NLQs to SQL using rule-based methods and custom heuristics, but extending these techniques to new databases and languages demands substantial manual effort and struggles with edge cases (Tang et al., 2025). Neural methods improve portability by casting text-to-SQL as sequence generation with integrated validation. PICARD (Scholak et al., 2021) uses sequence-to-sequence models with modular syntactic checks to ensure query correctness and alignment with the NLQ. Supplying schema context further improves semantic fidelity (Hong et al., 2024). In practice, these ideas combine constrained decoding for syntax with schema-aware prompting or retrieval for semantics, directly informing our KQL setting. Unlike prior systems that center on a single model, our approach uses both SLMs and LLMs. These general principles have been extended to domain-specific query languages, especially in security, where KQL is the primary language for threat hunting and log analysis. NL2KQL (Tang et al., 2025) translates NLQs to KQL by combining embedding-based semantic similarity with few-shot prompting in a modular, end-to-end framework. XPert (Jiang et al., 2024) takes an incident-centric approach, retrieving similar historical tickets to recommend or compose queries using in-context learning with embedding-based retrieval. However, both systems rely heavily on LLMs, resulting in cost and latency constraints that limit real-time SOC deployment, where analysts need rapid query assistance. In contrast, our approach uses an SLM for generation, avoiding exclusive dependence on LLMs.
@@
Security-Specific LLMs LLMs are increasingly applied across SOC tasks and offensive evaluations, but their effectiveness hinges on task-specific scaffolding and verification. VulDetect (Omar and Shiaeles, 2023) fine-tunes GPT-style models for anomaly detection in logs using a transformer-based framework. PentestGPT (Deng et al., 2024) builds an LLM penetration-testing assistant with modular task support, reporting bounded gains on real targets, highlighting the importance of structured tooling around the model. Although LLMs offer strong code generation capabilities, their cost and latency often make real-time deployment impractical (Belcak et al., 2025), pushing real-time operations toward SLMs.
@@
SLMs Adoption SLMs provide attractive latency and cost profiles for SOC settings but require augmentation to match LLM quality. SuperICL shows that small, locally fine-tuned models can act as plug-ins to larger LLMs, structuring context while the larger model executes—a useful division of labor for SOC workflows (Xu et al., 2024a). Parameter-efficient fine-tuning further narrows this gap while preserving efficiency; methods like LoRA and QLoRA enable targeted adaptation on modest hardware (Hu et al., 2022; Lin et al., 2025; Dettmers et al., 2023). Complementary techniques such as schema/value retrieval, few-shot selection, and best-of-N with an external judge improve correctness without enlarging model footprint (Wang et al., 2022; Madaan et al., 2023). In sum, SLMs are viable for SOC use when paired with retrieval, schema hints, and parameter-efficient tuning.
@@ 3.2. Zero-Shot Prompting
The Zero-Shot prompting strategy aims to interact with the model without containing any relevant examples or demonstrations (Saravia, 2024). Zero-Shot prompting can be an extremely powerful technique, as it typically utilizes less tokens than other prompting strategies and relies mostly on the inference capabilities of the language model itself. Furthermore, instruction tuning has been shown to improve Zero-Shot learning capabilities (Payong, 2024).
@@ 3.3. NL2KQL-Inspired Prompting
In addition to Zero-Shot prompting, we re-implement NL2KQL (Tang et al., 2025), the state-of-the-art system for KQL query generation. NL2KQL consists of five main components: Semantic Data Catalog, Schema Refiner, Few-Shot Selector, Prompt Builder, and Query Refiner. For detailed descriptions of these components, we refer the reader to the original paper (Tang et al., 2025). Below, we describe the changes we made in our recreation.
@@ 3.3.2. Schema Refiner
Consistent with (Tang et al., 2025), the Schema Refiner retrieves the top-$t$ relevant tables and associated columns using cosine similarity between the NLQ and the embeddings. As in the original, $t=9$ and $v_{\text{n}}=5$. We follow the same procedure but base our embeddings on the recreated Semantic Data Catalog.
@@ 3.3.4. Prompt Builder
We preserve the design of the Prompt Builder from (Tang et al., 2025), which combines detailed instructions, syntax rules, best practices, and few-shot examples into a single prompt template.
@@ 3.3.5. Query Refiner
We implement the Query Refiner following (Tang et al., 2025), which uses Microsoft’s KQL Parser (17) to verify syntactic and semantic correctness, detect undefined identifiers, repair aggregate functions, fix parentheses, and add missing operators. We retain the embedding-based replacement strategy with a cosine similarity threshold of 0.9.
@@ 3.3.6. Alternative Prompting Strategies
Beyond faithfully reimplementing NL2KQL (Tang et al., 2025), we introduce prompting strategies that supply models with targeted tips on common errors. The first focuses on frequent syntactic issues identified via the KQL Parser. The second adds guidance for semantic errors such as mismatched column types and invalid table references. We compare both strategies and adopt the one with the highest correctness. Due to SLMs’ limited inference capabilities, concise, model-specific tips help them generate more accurate KQL compared to supplying the full syntax and semantic rules. Our reimplementation substitutes Google’s text-embedding-004 for embeddings, uses Google Gemini 2.0 Flash and Microsoft Phi-4 for query generation, and shifts correctness filtering to post-generation. We further extend prompting with error-aware instructions to assess whether targeted guidance improves KQL generation.
@@ 3.4. LoRA Fine-Tuning with LLM Distillation
To further assess the quality of KQL queries that are generated from SLMs, we also test LoRA, a fine-tuning method that freezes pre-trained model weights from language models, and applies decompositional matrices to transformers layers within the LLM in order to alter a subset of model weights (Hu et al., 2022). The goal of this method is to train SLMs on NLQ-KQL pairs so that SLMs can be further improved in producing quality responses. LoRA can be mathematically defined as follows:
@@
Where B $\in$ $\mathbb{R}^{d\times r}$, A $\in$ $\mathbb{R}^{r\times k}$, and both are low-rank matrices. During training time, the values of the original weight matrix $\mathcal{W}$ remain stable while $\Delta{W}$ (ABAB) is updated. Altering low-rank matrices instead of the entire weight matrix reduces the memory requirements, gradient storage, and optimizer states needed to perform fine-tuning (Lin et al., 2025). To train the DeepSeek Coder 6.7B Instruct, we use a synthesized dataset created through LLM Knowledge Distillation.
LLM Knowledge Distillation is a transfer learning technique in Machine Learning that is used to relay the inference and reasoning capabilities of larger, proprietary LLMs to smaller SLMs. This allows the SLMs to preserve its nature of greater efficiency, deployment feasibility, and faster inference capabilities while also receiving the knowledge of LLMs in its own reasoning processes (Xu et al., 2024b).
@@ 3.5. Two-Staged Architecture
The oracle operates in two modes. In refinement-only, it selects and edits the best SLM candidate using the judge model’s internal knowledge. In schema-aware refinement, it performs the same process but with explicit schema context. We evaluate both modes using two prompts (Figure 2). The oracle is instantiated as an LLM-as-a-Judge, leveraging evidence that judge models align well with human preferences while exhibiting recognizable biases (Zheng et al., 2023).
@@ 4. Evaluation
We evaluate the effectiveness of our proposed design through a series of experiments conducted on a machine with an AMD EPYC 9534 64-Core Processor, an NVIDIA H100 NVL GPU, and Ubuntu 22.04.5 LTS. We address seven research questions (RQs) as follows. For all RQs, we access and query SLMs using the Huggingface Python package (Hugging Face, 2016), which avoids input/output token rate limits. For RQ5 and RQ6, we query Google Gemini 2.0 Flash via the Google GenAI package (Google, 2023). Our re-implemented NL2KQL pipeline (Section 3.3) is evaluated using Gemini 2.0 Flash, Phi-4, and GPT-4o.
@@
Datasets RQ1-RQ5 are evaluated using Microsoft’s NL2KQL Defender Evaluation Dataset (22), which contains 230 NLQ-KQL pairs referencing all tables in the NL2KQL Defender Schema. Instead of feeding NLQs directly to the models, we embed them in different prompting strategies and compare the generated queries against the baseline pairs to measure syntactic, semantic, and structural similarity. For RQ6, we obtained 83 NLQ-KQL pairs from the Microsoft Defender sections of Sentinel-Queries (Zorich, 2021), a publicly available GitHub repository. We used all available Defender queries from this repository and excluded queries designed for other data sources.
Evaluation Metrics In order to evaluate these research questions and assess the effectiveness of our research methods, we first define the following evaluation metrics. These evaluation metrics have been previously defined in NL2KQL (Tang et al., 2025), and we adapt them as part of our evaluation process as well.
@@
Cost:. Total cost (USD) to process 230 NLQs from the NL2KQL Defender Evaluation Dataset, accounting for all input and output tokens per model. SLMs offer several advantages, including enhanced data security and lower response latency, but cost efficiency is one of their most significant benefits. Like LLMs, SLM costs are determined by input and output token counts, with per-million-token pricing varying by provider. For this analysis, we report prices based on Cloudflare’s Workers AI pricing, as highlighted in (1). Table 1 presents the cost per 1M tokens for multiple LLMs and SLMs in USD. The cost gap is substantial: SLMs can offer up to 7.5x savings on input tokens alone compared to LLMs. When used effectively for code generation, SLMs can deliver significant performance at a fraction of traditional LLM costs.
@@
Embedding Model An embedding model is an ML model that can be used to convert text into a numerical representation so that it can be conceptually understood by LLMs or SLMs. The primary embedding model used in the following RQs is Google’s text-embedding-004 model. The text-embedding-004 provides embedding queries for \$0.025 per 1 million input tokens, and \$0.02 per 1 million output tokens (Google, 2025). Furthermore, text-embedding-004 is a lightweight, simplistic embedding model that has performed well across multiple numerous benchmarks and has strong multilingual and domain-specific capabilities.